9. Graph the line passing through (-2,-4) with a slope of Ys. 6 5 4 2 1 -6-5-4-3-2 3 4 5 6 2 3 torte.com Jounismo ofio
Answer #1

g ) let (2o, 7o)= (-2,-4) fm = 1/3 .. Equation of line is Y-70 = m Cae-360) y-6-9 = (32-627) yth= 13 (x+2) this is equation o